Juan is 90 years younger than Stacy. 2 years ago, Stacy's age was 4 times Juan's age. How old is Juan now?
Answer by ikleyn(52787)

.
Let J be the Juan'a age.


Then the Stacy age is (J+90), according to the condition.


2 years ago:


    (J+90)-2 = 4*(J-2)


    J + 88 = 4J - 8


    88 + 8 = 4J - J


    96 = 3J  ====>  J = 96/3 = 32.


Answer.  Juan is 32 years old now.